OpenStudy (meilendurcer):

Joy knits a square blanket that has an area of 1,500 square inches. What is the approximate length of each side of the blanket?

OpenStudy (anonymous):

Ok, now, squares have the same length and width. The formula to find area is length times width. So, x will represent our length, and if x is our length, then x would be our width, since they share the same length. So, x^2=1500, now,, solve :)

OpenStudy (anonymous):

what is left unexplained?

OpenStudy (anonymous):

square has a 4 equal sides

OpenStudy (meilendurcer):

i have no idea how to do this math okay

OpenStudy (anonymous):

Area=x^2

OpenStudy (anonymous):

where x is one of the sides of square

OpenStudy (anonymous):

1500=x^2 x=38.72